Solving Equations by the Addition Property
Solve and check the following equations
5x + 8 + 3x – x + 5 = 6x - 3

Answer by checkley75(3666)

5x+8+3x-x+5=6x-3
5x+3x-x-6x=-8-5-3
x=-16 answer
proof
5*-16+8+3*-16-16+5=6*-16-3
-80+8-48+16+5=-96-3
-128+29=-99
-99=-99